40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 30 19:43:33 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching 
    Reply  Wed Mar 30 19:54:19 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching laser_output_non_circular.png
       Reply  Wed Mar 30 20:08:34 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching Beam_Matching_02c_Vertical.pngBeam_Matching_02c_Horizontal.png
          Reply  Wed Mar 30 20:14:17 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching 
             Reply  Wed Mar 30 20:31:30 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching 
                Reply  Wed Mar 30 20:50:43 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching After_Faraday.pngAfter_Faraday_and_Rotated_Lens.png
                   Reply  Wed Mar 30 21:21:15 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching Oven_Lens_Solution_66mm.pngMatching_Into_Green_Oven_zoomed_out.pngMatching_Into_Green_Oven_zoomed_in.png
                      Reply  Thu Mar 31 02:59:49 2011, Koji, Configuration, Green Locking, The wonderful world of mode-matching 
                         Reply  Thu Mar 31 14:10:00 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching P3310028.JPGP3310029.JPG
                         Reply  Thu Mar 31 15:23:14 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching Zoomed_out_fitting01.pngZoomed_in_fitting01.png
                            Reply  Mon Apr 4 14:20:32 2011, Bryan, Configuration, Green Locking, The wonderful world of mode-matching Waist_Fits_from_laser.pngWaist_Fits_Bench_Position.png
Message ID: 4476     Entry time: Thu Mar 31 14:10:00 2011     In reply to: 4473
Author: Bryan 
Type: Configuration 
Category: Green Locking 
Subject: The wonderful world of mode-matching 

Quote:

 I went through the entries.

1. Give us a photo of the day. i.e. Faraday, tilted lens, etc...

2. After all, where did you put the faraday in the plot of the entry 4466?

3. Zoomed-in plot for the SHG crystal show no astigmatism. However, the zoomed out plot shows some astigmatism.
How consistent are they? ==> Interested in seeing the fit including the zoomed out measurements.

 OK. Taking these completely out of order in the easiest first...

2. The FI is between positions 27.75 and 32 on the bench - i.e. this is where the input and output apertures are. (corresponds to between 0.58 and 0.46 on the scale of those two plotsand just before both the vertical and horizontal waists) At these points the beam radius is around 400um and below, and the aperture of the Faraday is 4.8mm (diameter).

1. Photos...

Laser set up - note the odd angles of the mirrors. This is where we're losing a goodly chunk of the light. If need be we could set it up with an extra mirror and send the light round a square to provide alignment control AND reduce optical power loss...

P3310028.JPG

 

Faraday and angled lens - note that the lens angle is close to 45 degrees. In principle this could be replaced with an appropriate cylindrical lens, but as long as there's enough light passing through to the oven I think we're OK.

P3310029.JPG

3. Fitting... coming soon once I work out what it's actually telling me. Though I hasten to point out that the latter points were taken with a different laser power setting and might well be larger than the actual beam width which would lead to astigmatic behaviour.

ELOG V3.1.3-